Find three consecutive integers such that the square of the smallest is 29 less than the product of the larger two.
Answer by ikleyn(52835)

.
n^2 = (n+1)*(n+2) - 29


n^2 = n^2 + n + 2n + 2 - 29


3n = 27  ====>  n = 9.


Answer.  9, 10 and 11.

Solved.